# Coupling quantum-like cognition with the neuronal networks within generalized probability theory

### v1

@@ 4.6 Interference effect
We remark that the interference of probabilities is one of the basic quantum effects; for example, this effect can be observed in the two slit experiment (as was emphasized by Feynman [34, 35], see also [48]). In cognitive science and decision making it was analyzed in [50] as a nonclassical property of cognitive information processing, its coupling with the disjunction effect in cognitive psychology was established in [19]. We remark that in classical probability theory probabilities don’t interfere.
@@ 6 Model 1: Compound systems, tensor product representation
The graph of the $S_{1}\otimes S_{2}$ description can be constructed as follows. Denote the edges of the directed graphs $G_{S_{p}},p=1,2$, by the symbols $e^{(p)}_{ij}$. These are the vertexes of the graph under construction that is symbolically denoted as $G_{S_{1}\otimes S_{2}}$. This graph is undirected and weighted, and any vertex $e^{(1)}_{ij}$ is connected with any vertex $e^{(2)}_{km}$ by the edge $e^{(12)}_{ijkm}$ with the weight $w_{ijkm}$. For example, for $N_{1}=2,N_{2}=3$, the graph $G_{S_{1}\otimes S_{2}}$ has $32$ (undirected) edges and the graph $G_{S_{1}\cup S_{2}}$ has 25 (directed) edges. The graph $G_{S_{1}\otimes S_{2}}$ is not the standard tensor product $G_{S_{1}}\otimes G_{S_{2}}$ of the graphs $G_{S_{1}}$ and $G_{S_{2}}$. The latter graph will be used for the description of compound systems in models 2, 3 (see appendix). The graph $G_{S_{1}\otimes S_{2}}$ can be called the edge tensor product of the directed graphs $G_{S_{1}}$ and $G_{S_{1}}$.
@@ 7 Concluding remarks
As was pointed out, quantum-like modeling is actively developed as a phenomenological approach. But it really cries for coupling with brain’s functioning, with activity of the networks of communicating neurons. In this article the basics of such coupling were established within the GPT model based on the representation of networks by ordered weighted graphs that in turn are described by the matrices of weights for intensities of signaling between neurons. The graph-matrix GPT model (Model 1) has classical state space, a polygon, and its quantum-like properties are generated by the calculus of measurement instruments.
We presented some quantum-like effects, the order, interference (disjunction), and non-repeatability effects, within the graph-matrix GPT model. The compound systems, correlated neuronal networks, are considered only briefly and we plan to study quantum-like properties related to entanglement of the states of neuronal networks in one of future publications.
@@ Acknowledgments
A.K. was partially supported by the EU-grant CA21169 (DYNALIFE), by JST, CREST Grant Number JPMJCR23P4, Japan, and visiting professor fellowship to Tokyo University of Science (April 2024). M.O. was partially supported by JSPS KAKENHI Grant Numbers JP24H01566, JP22K03424, JST CREST Grant Number JPMJCR23P4, RIKEN TRIP Initiative (RIKEN Quantum), and the Quantinuum–Chubu University Collaboration in 2023–2024.

### v2

@@ 4.6 Interference effect
We note that probability interference is one of the fundamental quantum effects. For example, this effect can be observed in the double-slit experiment, as emphasized by Feynman [34, 35], and further discussed in [48]. In cognitive science and decision-making, it was analyzed in [50] as a nonclassical property of cognitive information processing. Its connection with the disjunction effect in cognitive psychology was established in [19]. It is important to note that in classical probability theory, probabilities do not interfere.
@@ 6 Model 1: Compound systems, tensor product representation
The graph of the $S_{1}\otimes S_{2}$ description can be constructed as follows. Denote the edges of the directed graphs $G_{S_{p}},p=1,2$, by the symbols $e^{(p)}_{ij}$. These are the vertexes of the graph under construction that is symbolically denoted as $G_{S_{1}\otimes S_{2}}$. This graph is undirected and weighted, and any vertex $e^{(1)}_{ij}$ is connected with any vertex $e^{(2)}_{km}$ by the edge $e^{(12)}_{ijkm}$ with the weight $w_{ijkm}$. For example, for $N_{1}=2,N_{2}=3$, the graph $G_{S_{1}\otimes S_{2}}$ has $32$ (undirected) edges and the graph $G_{S_{1}\cup S_{2}}$ has 25 (directed) edges. The graph $G_{S_{1}\otimes S_{2}}$ is not the standard tensor product $G_{S_{1}}\otimes G_{S_{2}}$ of the graphs $G_{S_{1}}$ and $G_{S_{2}}$. The latter graph will be used for the description of compound systems in models 2, 3. The graph $G_{S_{1}\otimes S_{2}}$ can be called the edge tensor product of the directed graphs $G_{S_{1}}$ and $G_{S_{1}}$. Since in such a graph the structure of the original neuronal graphs $S_{1}$ and $S_{2}$ would be shadowed, we illustrate the tensor product description of the compound system $S=(S_{1},S_{2})$ by the neuronal graphs for $S_{1}$ and $S_{2}$ and correlation coupling between them, so see Fig. 2 for such vizualization
0,1; Refer to caption
Figure 2: Visualization of correlations between signaling within two networks $S_{1}$ and $S_{2}$ (top panel) and the direct physical connection between them (bottom panel), see Remark for further clarification and discussion.
Remark. Each of two networks $S_{1}$ and $S_{2}$ is characterized by internal signaling, described by the random signals $\xi_{ij}$ in $S_{1}$ and $\xi_{ij}^{\prime}$ in $S_{2}$ (for simplicity of visualization some possible connections between neurons are omitted). This is the direct signaling within the networks. Besides this direct signaling, there exist temporal concurrences between the signals in $S_{1}$ and $S_{2}$ (propagating along the edges in each of them). Mathematically they can be represented as correlations between the random signals $\xi_{ij}$ and $\xi_{ij}^{\prime},$ see (38). These correlations are basic for the quantum-like representation via the tensor product. Of course, there can exist direct physical connections between the neurons from $S_{1}$ and $S_{2}$ (see bottom panel of Fig. 2). In the tensor product representation such direct signaling between the neurons from these two networks are ignored. The latter can be justified only if the number of such direct $S_{1}\leftrightarrow S_{2}$ connections is very small compare to the numbers of internal connections within $S_{1}$ and $S_{2}.$ Otherwise the correlation matrix would not reflect the functioning of these two networks. Hence, “networks entanglement” physically means very dense connectivity (intensive signaling) within each of them in combination with poor inter-network connectivity (weak inter-network signaling), cf. with [83]. The crucial question is to understand how the compound network can explore the correlations between signals within them - in the absence of physical connections. We can only speculate that the electromagnetic field within the brain should be involved in establishing of such inter-network and generally long-distance correlations.
@@ 7 Concluding remarks
As noted, quantum-like modeling is actively developing as a phenomenological approach. However, it strongly requires integration with brain functioning, specifically with the activity of networks of communicating neurons. In this article, we established the basics of such integration within the GPT model, which represents networks using ordered weighted graphs. These graphs are described by weight matrices that quantify the intensities of signaling between neurons. The graph-matrix GPT model (Model 1) has a classical state space — a polygon — and its quantum-like properties are generated by the calculus of measurement instruments.
We presented several quantum-like effects — such as the order, interference (disjunction), and non-repeatability effects — within the graph-matrix GPT model. The discussion of compound systems and correlated neuronal networks was brief, and we plan to explore the quantum-like properties related to the entanglement of neuronal network states in a future publication.
@@ Acknowledgments
A.K. was partially supported by the EU-grant CA21169 (DYNALIFE), by JST, CREST Grant Number JPMJCR23P4, Japan, and visiting professor fellowship to Tokyo University of Science (April 2024). M.O. was partially supported by JSPS KAKENHI Grant Numbers JP24H01566, JP22K03424, JST CREST Grant Number JPMJCR23P4, RIKEN TRIP Initiative (RIKEN Quantum), and the Quantinuum–Chubu University Collaboration in 2023–2024. A.K. would like to thank prof. Bussmeyer for supporting discussion on the